OpenStudy (anonymous):

The ratio of Sue's age to Bob's age is 3 to 7. The ratio of Sue's age to Joe's is 4 to 9. What is the ratio of Bob's age to Joe's ?

OpenStudy (anonymous):

let Sue's age=a, Bob's age =b, Joe's =c \[\frac{a}{b}=\frac{3}{7}\] \[\frac{a}{c}=\frac{4}{9}\] we want b/c \[\frac{b}{c}=\frac{4}{9} * \frac{7}{3}\] ratio of Bob's age to Joe's =28/27